Two cars approach each other on a straight and level road. Car A is traveling at 75 km/h, due north and car B is traveling at 45 km/h, due south. Both velocities are measured relative to the ground. What is the velocity of car A relative to an observer in car B?


1
Expert's answer
2020-09-07T08:55:48-0400

The velocity of car A relative to an observer in car B is the same as velocity of B relative to A:


"v=v_A+v_B=75+45=120\text{ km/h}."